# Parameter-Efficient Fine-Tuning of LLaMA for the Clinical Domain

### v1

@@ Abstract
Adapting pretrained language models to novel domains, such as clinical applications, traditionally involves retraining their entire set of parameters. However, this approach is increasingly proven to be impractical owing to the substantial computational requirements associated with training such large language models. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) techniques offer a viable solution by selectively fine-tuning a small subset of additional parameters, significantly reducing the computational requirements for domain adaptation. In this study, we propose Clinical LLaMA-LoRA, a PEFT adapter layer built upon the open-sourced LLaMA model. Clinical LLaMA-LoRA is trained using clinical notes obtained from the MIMIC-IV database, thereby creating a specialised adapter designed for the clinical domain. Additionally, we propose a two-step PEFT framework which fuses Clinical LLaMA-LoRA with Downstream LLaMA-LoRA, another PEFT adapter specialised for downstream tasks. We evaluate this framework on multiple clinical outcome prediction datasets, comparing it to clinically trained language models. Our proposed framework achieves a state-of-the-art AUROC score averaged across all clinical downstream tasks. We observe substantial improvements of 6-9% AUROC score in the large-scale multilabel classification tasks, such as diagnoses and procedures classification.
@@ 1 Introduction
In this study, we propose a two-step PEFT framework (see Figure 1). Firstly, we introduce Clinical LLaMA-LoRA, a Low-Rank Adaptation (LoRA, Hu et al., 2022) PEFT adapter built upon the open-source Large Language Model Meta AI (LLaMA) Touvron et al. (2023). Then, we introduce Downstream LLaMA-LoRA, which is trained on top of the pretrained Clinical LLaMA-LoRA. Downstream LLaMA-LoRA is specifically designed for clinical downstream tasks. The fusion of the two adapters achieves state-of-the-art performance in clinical NLP downstream tasks while considerably reducing the computational requirements. This study presents the following contributions:
@@ 3.1 Problem Statement
$\operatorname*{argmax}_{\theta}\sum_{(x,y)\in\mathcal{Z}}\sum_{t=1}^{|y|}\log\left(P_{\Phi+\Delta\Phi(\theta)}\left(y_{t}\mid x,y_{<t}\right)\right)$
@@
3.2 Clinical LLaMA-LoRA
Clinical LLaMA-LoRA is a LoRA adapter built upon LLaMA Touvron et al. (2023). Clinical LLaMA-LoRA is domain-adapted to the clinical domain and fine-tuned to the downstream tasks following the proposed procedure shown on the right-hand side of Figure 2.
@@ 3.2 Clinical LLaMA-LoRA
Domain-adaptive Pretraining
Clinical LLaMA-LoRA is trained using a combination of MIMIC-IV de-identified discharge summaries (331,794) and radiology reports (2,321,355), resulting in a collection of 2,653,149 individual clinical notes. We evaluate five different PEFT techniques, which include LoRA, Adaptation Prompt, Prefix Tuning, Prompt Tuning, and P-tuning.
@@
Downstream Fine-tuning
@@
  • Prolonged mechanical ventilation (PMV): a binary classification task to predict whether a patient will require mechanical ventilation for more than seven days Huang et al. (2020).
  • In-hospital mortality (MOR): a binary classification task to predict whether a patient will survive during their hospital stay van Aken et al. (2021).
  • Length of stay (LOS): a multiclass classification task to predict the length of a patient’s hospital stay, categorised into four time-bins: less than three days, three to seven days, one to two weeks, and more than two weeks van Aken et al. (2021).
  • Diagnoses (DIAG): a large-scale multilabel classification task to predict the differential diagnoses associated with a patient, represented by simplified ICD-9 diagnosis codes van Aken et al. (2021).
  • Procedures (PROC): a large-scale multilabel classification task to predict the diagnostics or treatments administered to a patient, represented by simplified ICD-9 procedure codes van Aken et al. (2021).
@@ 3.3 Baseline Models
The baseline models used in the evaluation are as follows:
  • Bio+ClinicalBERT Alsentzer et al. (2019): Bio+ClinicalBERT is pretrained on clinical notes from the MIMIC-III database. It is initialised from a biomedical language model called BioBERT Lee et al. (2019), which is pretrained on biomedical research articles.
  • BlueBERT Peng et al. (2019): BlueBERT is pretrained on clinical notes from the MIMIC-III database and PubMed abstracts starting from the pretrained checkpoint of BERT Devlin et al. (2019), a general-domain language model.
  • CORe van Aken et al. (2021): CORe is pretrained on clinical notes from the MIMIC-III database and biomedical articles starting from the pretrained checkpoint of BioBERT Lee et al. (2019).
  • UmlsBERT Michalopoulos et al. (2021): UmlsBERT is pretrained on clinical notes from the MIMIC-III database starting from the pretrained checkpoint of Bio+ClinicalBERT while modifying the architecture and pretraining objective by incorporating knowledge from the Unified Medical Language System (UMLS) Metathesaurus Schuyler et al. (1993).
These baseline models have been trained to perform specifically on clinical data, thus providing comparison points for evaluating the performance of the proposed Clinical LLaMA-LoRA in downstream clinical NLP tasks.
@@ 4 Results and Analysis
4.1 Pretraining
Base Model PEFT Trainable Params Train Perplexity Test Perplexity Train Time (h:m:s)
LLaMA LoRA 8,388,608 (0.12%) 1.858 2.244 21:37:42
Adaptation Prompt 1,228,830 (0.02%) 2.561 2.865 24:57:17
Prefix Tuning 5,242,880 (0.08%) 2.815 2.748 20:11:07
Prompt Tuning 61,440 (0.0009%) 4.846 4.007 23:27:28
P-tuning 16,093,696 (0.24%) 2.723 3.271 23:49:31
PMC-LLaMA LoRA 2,097,152 (0.03%) 1.938 2.404 21:32:59
Adaptation Prompt 1,228,830 (0.018%) 2.374 2.867 23:33:10
Prefix Tuning 2,621,440 (0.04%) 1.789 2.848 20:13:10
Prompt Tuning 40,960 (0.0006%) 4.821 4.385 22:25:32
P-tuning 2,171,392 (0.03%) 3.491 4.572 22:28:15
@@ 4.1 Pretraining
The pretraining results can be found in Table 2. We employ PEFT techniques to perform domain-adaptive pretraining. All PEFT techniques train a significantly smaller number of parameters, ranging from only 0.001% to 0.24% of the original model parameters, which substantially decreases the computational resources required and shortens the training time. Note that performing full-parameter training of LLaMA and PMC-LLaMA with just a single GPU is unfeasible. Instead, PEFT techniques require less than 24 hours per epoch on average with only a single NVIDIA A100-80GB GPU.
Among all the PEFT techniques, LoRA emerges as the best-performing one for both LLaMA and PMC-LLaMA in the clinical domain-adaptive pretraining, achieving the lowest perplexity scores of 2.244 and 2.404, respectively. This pretrained LoRA is referred to as Clinical LLaMA-LoRA in the subsequent sections. The following experiments in downstream fine-tuning will utilise this pretrained Clinical LLaMA-LoRA.
4.2 Downstream results
@@
BlueBERT 53.12 76.95 66.36 73.81 77.70 69.59
UmlsBERT 55.49 75.87 66.06 64.34 74.19 67.19
Bio+ClinicalBERT 54.49 72.92 65.13 65.97 71.73 66.05
CORe 52.11 71.52 64.17 72.40 72.73 66.59
@@
     + Clinical LLaMA-LoRA (Frozen) 51.62 65.66 58.16 63.47 69.01 61.58
         + Downstream LLaMA-LoRA 51.11 66.00 58.04 60.46 65.30 60.18
     + Clinical LLaMA-LoRA (Trainable) 55.76 74.81 64.83 76.07 82.76 70.85
         + Downstream LLaMA-LoRA 56.72 76.99 65.86 78.29 86.17 72.81
@@
Table 3: AUROC scores in clinical downstream document classification tasks. The macro-averaged AUROC score is calculated by taking the average of AUROC scores across all tasks. The boldface cell indicates the highest AUROC score in a column, the row in italic indicates the model variant with the highest macro-averaged AUROC in its category. $\ast$ Due to restricted computing resources, the fine-tuning of LLaMA and PMC-LLaMA was constrained to only training the final classification layer.
Model PMV MOR LOS DIAG PROC Macro Average
BlueBERT 53.12 76.95 66.36 73.81 77.70 69.59
     + LoRA 55.77 81.90 70.48 70.66 78.10 71.56
UmlsBERT 55.49 75.87 66.06 64.34 74.19 67.19
     + LoRA 56.59 80.33 69.03 69.68 77.53 70.63
BioClinicalBERT 54.49 72.92 65.13 65.97 71.73 66.05
     + LoRA 56.13 78.81 68.28 68.53 75.19 69.39
CORe 52.11 71.52 64.17 72.40 72.73 66.59
     + LoRA 55.31 79.27 68.18 67.34 72.36 68.49
LLaMA + Clinical LLaMA-LoRA + Downstream LoRA 56.72 76.62 65.86 78.29 86.17 72.81
Table 4: AUROC scores of the LoRA-equipped baseline models in clinical downstream tasks. The boldface cell indicates the highest AUROC score in a column. The row in italic indicates the model variant with the highest macro-averaged AUROC in its category.
@@
We compare the results obtained by LLaMA and LLaMA + LoRA, as well as PMC-LLaMA and PMC-LLaMA + LoRA, as presented in Table 3. The obtained results consistently demonstrate improved AUROC scores when utilising LoRA across all tasks. The macro-averaged AUROC score of LoRA-equipped LLaMA shows a notable 13.01% increase when compared to the LLaMA-only baseline. Similarly, LoRA-equipped PMC-LLaMA exhibits a 12.2% improvement in macro-averaged AUROC compared to the original PMC-LLaMA Both LLaMA and PMC-LLaMA, when equipped with LoRA, exhibit significant AUROC score improvements in all tasks except the prolonged mechanical ventilation prediction task, which is proven challenging for all model variants.
Furthermore, the marginal difference in AUROC scores between PMC-LLaMA and the general-domain LLaMA can be attributed to two factors. Firstly, the original LLaMA has been exposed to biomedical concepts during its pretraining, reducing the need for domain-adaptive pretraining to the biomedical domain. Secondly, clinical NLP tasks are challenging, even for biomedical LLMs.
@@
We compare the AUROC scores obtained by the baseline models, and LoRA-equipped LLaMA and PMC-LLaMA (see Table 3). Among the baseline models, BlueBERT performs the best with a macro-averaged AUROC score of 69.59%. Compared to BlueBERT, both LLaMA and PMC-LLaMA underperform with macro-averaged AUROC scores of 58.61% and 60.51%, respectively. This finding highlights the importance of clinical-specific fine-tuning.
Significant improvements can be observed in LoRA-equipped LLaMA and PMC-LLaMA, with macro-averaged AUROC scores of 71.62% and 72.71%, respectively. We notice considerable improvements in the diagnoses and procedures prediction tasks. For example, LoRA-equipped LLaMA achieves AUROC scores of 78.37% and 87.49% in the diagnoses and procedures prediction tasks, respectively, compared to 73.81% and 77.70% for BlueBERT. This represents improvements of 4.56% in diagnoses prediction and 9.79% in procedures prediction. Improvements are also observed in the results obtained by LoRA-equipped PMC-LLaMA, outperforming BlueBERT by 5% in diagnoses prediction and 9.02% in procedures prediction.
Overall, LoRA-equipped LLaMA and PMC-LLaMA achieve higher AUROC scores than the baseline clinical LMs in various clinical prediction tasks, particularly in diagnoses, procedures, and mortality predictions, while maintaining competitive AUROC scores in length-of-stay prediction. However, LoRA-equipped LLaMA and PMC-LLaMA still underperform in prolonged mechanical ventilation prediction.
@@
The domain-adaptive pretraining step yields the clinically-trained LoRA adapters for LLaMA and PMC-LLaMA, called Clinical LLaMA-LoRA. We compare the results of Clinical LLaMA-LoRA-equipped LLaMA and PMC-LLaMA with the baseline models. We evaluate Clinical LLaMA-LoRA with and without downstream fine-tuning, referred to as "Trainable" and "Frozen" respectively.
The results indicate that Clinical LLaMA-LoRA-equipped LLaMA and PMC-LLaMA outperform the baseline models. LLaMA with a trainable Clinical LLaMA-LoRA achieves an AUROC score of 70.85%, surpassing BlueBERT’s score of 69.59%. PMC-LLaMA with a trainable Clinical LLaMA-LoRA achieves an even higher AUROC score of 72.23%. These findings demonstrate that the Clinical LLaMA-LoRA contributes to higher AUROC scores for LLaMA and PMC-LLaMA over clinically trained LLMs.
@@
Firstly, we evaluate the frozen pretrained Clinical LLaMA-LoRA. Both LLaMA and PMC-LLaMA with frozen Clinical LLaMA-LoRA do not exhibit a significant increase in performance compared to the original fine-tuning. This indicates that, despite the domain-adaptive pretraining, the limited number of trainable parameters during the downstream fine-tuning restricts the potential improvement that the model can achieve.
This reasoning is further supported by the significant improvement observed in the AUROC scores of LLaMA and PMC-LLaMA with trainable Clinical LLaMA-LoRA. LLaMA and PMC-LLaMA with trainable Clinical LLaMA-LoRA achieve 70.85% and 72.23% macro-averaged AUROC scores, respectively, massive improvements from the vanilla fine-tuning performance (58.61% and 60.51% AUROC scores respectively).
However, Clinical LLaMA-LoRA does not yield significant improvements when compared to LLaMA and PMC-LLaMA, which are directly equipped with LoRA without pretraining. For instance, we can observe that LLaMA with LoRA achieves a slightly higher macro-averaged AUROC score of 71.62% compared to LLaMA with Clinical LLaMA-LoRA, which achieves 70.85%.
@@
By considering Clinical LLaMA-LoRA as the "delta-updating" outcome of the domain-adaptive pretraining, we can view the downstream fine-tuning process as an additional "delta-updating" step. To investigate the impact of this approach, we conduct experiments by adding a Downstream LLaMA-LoRA to LLaMA and PMC-LLaMA models that were already equipped with Clinical LLaMA-LoRA. From Table 3, we can observe that Downstream LLaMA-LoRA fails to improve the performance of LLaMA and PMC-LLaMA with frozen Clinical LLaMA-LoRA. On the other hand, improvement can be observed when adding Downstream LLaMA-LoRA to LLaMA with trainable Clinical LLaMA-LoRA. This combination of LLaMA with trainable Clinical LLaMA-LoRA and Downstream LLaMA-LoRA achieves the highest macro-averaged AUROC score of 72.81%. The macro-averaged AUROC score of Clinical LLaMA-LoRA was almost similar to that of PMC-LLaMA with LoRA, suggesting similar efficacy between Clinical LLaMA-LoRA and the full fine-tuning process that PMC-LLaMA has undergone. Moreover, Clinical LLaMA-LoRA offers the advantage of reduced computational resources and training time, which is aligned with the requirements of practical implementation in clinical settings.
Can LoRA help better fine-tune clinically-trained LMs?
The baseline models are relatively smaller in size compared to the LLaMA-based models, which may be a better fit to care providers with limited access to computing resources. To that end, we experimented with fine-tuning the baseline models with LoRA.
Table 4 shows the obtained results. All baseline models see improvements in AUROC scores in all tasks. For instance, the LoRA-equipped BlueBERT achieves an improved macro-averaged AUROC score of 71.56% compared to the conventional fine-tuning with 69.59%.
This finding highlights the possibility of using LoRA to efficiently fine-tune clinically trained LMs, such as BlueBERT, to downstream use cases.

### v3

@@ Abstract
Adapting pretrained language models to novel domains, such as clinical applications, traditionally involves retraining their entire set of parameters. Parameter-Efficient Fine-Tuning (PEFT) techniques for fine-tuning language models significantly reduce computational requirements by selectively fine-tuning small subsets of parameters. In this study, we propose a two-step PEFT framework and evaluate it in the clinical domain. Our approach combines a specialised PEFT adapter layer designed for clinical domain adaptation with another adapter specialised for downstream tasks. We evaluate the framework on multiple clinical outcome prediction datasets, comparing it to clinically trained language models. Our framework achieves a better AUROC score averaged across all clinical downstream tasks compared to clinical language models. In particular, we observe large improvements of 4-5% AUROC in large-scale multilabel classification tasks, such as diagnoses and procedures classification. To our knowledge, this study is the first to provide an extensive empirical analysis of the interplay between PEFT techniques and domain adaptation in an important real-world domain of clinical applications. [^1] The code is accessible via https://github.com/aryopg/clinical_peft.
Parameter-Efficient Fine-Tuning of LLaMA for the Clinical Domain
Aryo Pradipta Gema1  Pasquale Minervini1  Luke Daines2 Tom Hope3,4Beatrice Alex5,6 1School of Informatics, University of Edinburgh  2Usher Institute, University of Edinburgh 3Allen Institute of AI 4Hebrew University of Jerusalem 5Edinburgh Futures Institute, University of Edinburgh 6School of Literatures, Languages and Cultures, University of Edinburgh {aryo.gema, p.minervini, luke.daines, b.alex}@ed.ac.uk tomh@allenai.org
@@ 1 Introduction
In this study, we propose a two-step PEFT framework (see Figure 1). Firstly, we introduce Clinical LLaMA-LoRA, a Low-Rank Adaptation (LoRA, Hu et al., 2022) PEFT adapter built upon the open-source Large Language Model Meta AI (LLaMA) Touvron et al. (2023). Then, we introduce Downstream LLaMA-LoRA, which is trained on top of the pretrained Clinical LLaMA-LoRA. Downstream LLaMA-LoRA is specifically designed for clinical downstream tasks. The fusion of the two adapters achieves better performance in clinical NLP downstream tasks compared to clinically trained LLMs while considerably reducing the computational requirements. This study presents the following contributions:
@@ 2.3 Parameter-Efficient Fine-Tuning for Large Language Models
2.4 Multi-step Adaptation
Prior studies have explored the two-step adaptation framework, although they have fundamental differences from our proposed setup. For instance, Zhang et al. (2021) introduced a multi-domain unsupervised domain adaptation (UDA) with a two-step strategy, involving domain-fusion training with Masked Language Model loss on a mixed corpus, followed by task fine-tuning with a task-specific loss on the domain corpus. More recently, Malik et al. (2023) introduced UDApter which utilises PEFT adapters to do efficient UDA. However, unsupervised domain matching techniques such as UDApter rely on restrictive assumptions about the underlying data distributions that are often unsatisfied in real-world scenarios Li et al. (2020). In our study, we experiment with the clinical domain as the target domain that is not available in the LLM’s initial pretraining. Consequently, significant discrepancies exist between the distributions of the source and target domains. Leveraging the amount of available clinical notes, we adopt a self-supervised learning paradigm by continually pretraining the LLMs within the target domain rather than relying on the UDA paradigm.
Our approach shares theoretical similarities with the multi-step continual pretraining approach, proposed by Gururangan et al. (2020), which proposes domain- and task-adaptive pretraining. However, the main difference between our proposed approach and Gururangan et al. (2020) is in the discrepancy between the source and the target domains. Gururangan et al. (2020) experimented with adapting general-domain LLMs to domains encountered during their initial pretraining, such as news and biomedical domains. On the other hand, we experiment with the clinical domain which is entirely absent from the LLMs’ initial pretraining due to legal constraints which restrict access to sensitive clinical notes. On top of that, adapting to the clinical domain poses a bigger challenge due to the complexity of medical terminology and the presence of incomplete sentences Lehman et al. (2023).
@@ 3.1 Problem Statement
$\operatorname*{argmax}_{\theta}\sum_{(x,y)\in\mathcal{Z}}\sum_{t=1}^{|y|}\log\left(P_{\Phi_{0}+\Delta\Phi(\theta)}\left(y_{t}\mid x,y_{<t}\right)\right)$
@@
3.2 Two-step LLaMA-LoRA
In this study, we propose a two-step PEFT framework (as shown on the right-hand side of Figure 2). Firstly, we introduce Clinical LLaMA-LoRA, a LoRA adapter built upon LLaMA Touvron et al. (2023) that is adapted to the clinical domain. Secondly, we introduce Downstream LLaMA-LoRA, which is trained on top of the pretrained Clinical LLaMA-LoRA and is specifically adapted to the downstream tasks.
@@ 3.2 Two-step LLaMA-LoRA
Domain-adaptive Pretraining: Clinical LLaMA-LoRA
Clinical LLaMA-LoRA is trained using a combination of MIMIC-IV de-identified discharge summaries (331,794) and radiology reports (2,321,355), resulting in a collection of 2,653,149 individual clinical notes. We evaluate five PEFT techniques, which include LoRA Hu et al. (2022), Adaptation Prompt Zhang et al. (2023), Prefix Tuning Li and Liang (2021), Prompt Tuning Lester et al. (2021), and P-tuning Liu et al. (2021b).
@@
Downstream Fine-tuning: Downstream LLaMA-LoRA
@@
  • Prolonged mechanical ventilation (PMV): a binary classification task to predict whether a patient will require mechanical ventilation for more than seven days Huang et al. (2020); Naik et al. (2022).
  • In-hospital mortality (MOR): a binary classification task to predict whether a patient will survive during their hospital stay van Aken et al. (2021); Naik et al. (2022).
  • Length of stay (LOS): a multiclass classification task to predict the length of a patient’s hospital stay, categorised into four time-bins: less than three days, three to seven days, one to two weeks, and more than two weeks van Aken et al. (2021); Naik et al. (2022).
  • Diagnoses (DIAG): a large-scale multilabel classification task to predict the differential diagnoses of a patient, represented by simplified ICD-9 diagnosis codes van Aken et al. (2021).
  • Procedures (PROC): a large-scale multilabel classification task to predict the treatments administered to a patient, represented by simplified ICD-9 procedure codes van Aken et al. (2021).
@@ 3.3 Baseline Models
We selected baseline models that have undergone a domain-adaptive pretraining process on clinical notes (MIMIC-III). Thus, these baseline models have been designed to perform specifically on clinical data, providing comparison points for evaluating our proposed approach of two-step adaptation in downstream clinical NLP tasks. The baseline models used in the evaluation are as follows:
  • Bio+ClinicalBERT Alsentzer et al. (2019): Bio+ClinicalBERT is pretrained on MIMIC-III clinical notes. It is initialised from a biomedical language model called BioBERT Lee et al. (2019), which is pretrained on biomedical research articles.
  • BlueBERT Peng et al. (2019): BlueBERT is pretrained on MIMIC-III clinical notes and PubMed abstracts starting from the pretrained checkpoint of BERT Devlin et al. (2019), a general-domain language model.
  • CORe van Aken et al. (2021): CORe is pretrained on MIMIC-III clinical notes and biomedical articles starting from the pretrained checkpoint of BioBERT Lee et al. (2019).
  • UmlsBERT Michalopoulos et al. (2021): UmlsBERT is pretrained on MIMIC-III clinical notes using the pretrained weights of Bio+ClinicalBERT with modified architecture and pretraining objective that incorporates knowledge from the Unified Medical Language System (UMLS) Metathesaurus Schuyler et al. (1993).
@@ 4 Results and Analysis
4.1 Domain-adaptive Pretraining
Base Model PEFT Trainable Params Train Ppl Test Ppl GPU Train Time (h:m:s)
Clinical LLaMA - 6.7B (100%) 1.811 2.210 4x80GB 49:26:38
LLaMA LoRA 8.4M (0.12%) 1.858 2.244 1x80GB 21:37:42
Adaptation Prompt 1.2M (0.02%) 2.561 2.865 1x80GB 24:57:17
Prefix Tuning 5.2M (0.08%) 2.815 2.748 1x80GB 20:11:07
Prompt Tuning 61.4K (0.0009%) 4.846 4.007 1x80GB 23:27:28
P-tuning 16.1M (0.24%) 2.723 3.271 1x80GB 23:49:31
PMC-LLaMA LoRA 2.1M (0.03%) 1.938 2.404 1x80GB 21:32:59
Adaptation Prompt 1.2M (0.018%) 2.374 2.867 1x80GB 23:33:10
Prefix Tuning 2.6M (0.04%) 1.789 2.848 1x80GB 20:13:10
Prompt Tuning 41K (0.0006%) 4.821 4.385 1x80GB 22:25:32
P-tuning 2.2M (0.03%) 3.491 4.572 1x80GB 22:28:15
@@ 4.1 Domain-adaptive Pretraining
The pretraining results can be found in Table 2. We employ PEFT techniques for domain-adaptive pretraining, requiring a significantly smaller number of parameters ranging from just 0.001% to 0.24% of the original model parameters. This approach substantially reduces the required computational resources and training time. We perform a full-parameter domain-adaptive pretraining of LLaMA, referred to as Clinical LLaMA, using four NVIDIA A100-80GB GPUs which took 49.5 hours. Instead, PEFT techniques require less than 24 hours per epoch on average with only a single GPU with a comparable perplexity score.
LoRA emerges as the best-performing PEFT method for both LLaMA and PMC-LLaMA in the clinical domain-adaptive pretraining, achieving the lowest perplexity scores of 2.244 and 2.404, respectively, which are very similar to Clinical LLaMA’s perplexity score of 2.210. This pretrained LoRA is referred to as Clinical LLaMA-LoRA in the subsequent sections. The following experiments in downstream fine-tuning will utilise this pretrained Clinical LLaMA-LoRA.
4.2 Downstream Fine-tuning
@@
BlueBERT 57.31 81.34 72.92 73.39 76.62 72.32
UmlsBERT 58.29 81.83 73.02 72.08 78.32 72.70
Bio+ClinicalBERT 54.00 72.67 72.21 76.65 83.21 71.75
CORe 52.11 71.52 64.17 72.40 84.51 69.40
Clinical LLaMA* 52.28 63.22 56.06 59.31 63.42 58.86
@@
       + Clinical LLaMA-LoRA (Frozen) 52.22 60.88 55.05 57.64 62.48 57.65
           + Downstream LLaMA-LoRA 52.31 61.72 55.16 57.70 62.58 57.90
       + Clinical LLaMA-LoRA (Trainable) 51.41 81.16 72.44 81.97 88.69 75.13
           + Downstream LLaMA-LoRA 53.81 83.02 73.26 81.93 88.31 76.07
@@
Table 3: AUROC scores in clinical downstream document classification tasks. The macro-averaged AUROC score is calculated by taking the average of AUROC scores across all tasks. The boldface cell indicates the highest AUROC score in a column, the row in italic indicates the variant with the highest macro-averaged AUROC in its category. + LoRA denotes applying LoRA on top of the pretrained LLM without domain-adaptive pretraining. + Clinical LLaMA-LoRA denotes applying Clinical LLaMA-LoRA that is domain-adaptively pretrained on top of the pretrained LLM. + Downstream LLaMA-LoRA denotes applying Downstream LLaMA-LoRA on top of the LLM + Clinical LLaMA-LoRA. Frozen means that the parameters are not trainable, while Trainable means that the parameters are trainable. $\ast$ Due to restricted computing resources, the fine-tunings of Clinical LLaMA, LLaMA, and PMC-LLaMA were constrained to only training the final classification layer.
@@
We compare the results obtained by LLaMA and LLaMA + LoRA, as well as PMC-LLaMA and PMC-LLaMA + LoRA, as presented in Table 3. The obtained results consistently demonstrate improved AUROC scores when utilising LoRA across all tasks. The macro-averaged AUROC score of LoRA-equipped LLaMA shows a notable 13.01% increase when compared to the LLaMA-only baseline. Similarly, LoRA-equipped PMC-LLaMA exhibits a 12.19% improvement in macro-averaged AUROC compared to the original PMC-LLaMA Both LLaMA and PMC-LLaMA, when equipped with LoRA, show significant AUROC score improvements in all tasks except the PMV prediction task, which is challenging for all model variants.
Furthermore, the marginal difference in AUROC scores between PMC-LLaMA and the general-domain LLaMA may be attributed to two factors. Firstly, the original LLaMA has been exposed to biomedical concepts during its pretraining, reducing the need for domain-adaptive pretraining to the biomedical domain. Secondly, clinical outcome prediction requires an understanding of how to apply biomedical knowledge in an interconnected manner to provide prognostic. We believe that biomedical pretraining may not be sufficient in providing such practical knowledge.
@@
We compare the AUROC scores obtained by the baseline models, and LoRA-equipped LLaMA and PMC-LLaMA (see Table 3). Among the baseline models, UmlsBERT performs the best with a macro-averaged AUROC score of 72.70%. Compared to UmlsBERT, both LLaMA and PMC-LLaMA underperform with macro-averaged AUROC scores of 58.61% and 60.51%, respectively. This finding highlights the importance of clinical-specific fine-tuning.
Significant improvements can be observed in LoRA-equipped LLaMA and PMC-LLaMA, with macro-averaged AUROC scores of 71.62% and 72.70%, respectively, with noticeable improvements in the diagnoses and procedures prediction tasks. LoRA-equipped LLaMA achieves AUROC scores of 78.37% and 87.49% in the diagnoses and procedures prediction tasks, respectively, compared to 72.08% and 78.32% for UmlsBERT. This represents improvements of 6.29% in diagnoses prediction and 9.17% in procedures prediction. Improvements are also observed in the results obtained by LoRA-equipped PMC-LLaMA, outperforming UmlsBERT by 6.73% in diagnoses prediction and 8.36% in procedures prediction.
@@
The domain-adaptive pretraining step yields the clinically-trained LoRA adapters for LLaMA and PMC-LLaMA, denoted as Clinical LLaMA-LoRA. We compare the results of Clinical LLaMA-LoRA-equipped LLaMA and PMC-LLaMA with the baseline models. We evaluate Clinical LLaMA-LoRA with and without fine-tuning, referred to as "Trainable" and "Frozen" respectively.
The results indicate that Clinical LLaMA-LoRA-equipped LLaMA and PMC-LLaMA outperform the baseline models. LLaMA with a trainable Clinical LLaMA-LoRA achieves an AUROC score of 75.13%, surpassing UmlsBERT’s score of 72.32%. PMC-LLaMA with a trainable Clinical LLaMA-LoRA achieves a lower AUROC score of 72.23%. LLaMA with a trainable Clinical LLaMA-LoRA also outperforms Clinical LLaMA which achieves an AUROC score of 58.86%.
These findings indicate that the Clinical LLaMA-LoRA contributes to higher AUROC scores for LLaMA and PMC-LLaMA over clinically trained LLMs, while biomedical domain-adaptive pretraining may not be necessary to improve the model’s performance in the clinical settings.
@@
Firstly, we evaluate the frozen pretrained Clinical LLaMA-LoRA. Both LLaMA and PMC-LLaMA with frozen Clinical LLaMA-LoRA do not exhibit a significant increase in performance compared to the original fine-tuning. This indicates that, despite the domain-adaptive pretraining, the limited number of trainable parameters during the downstream fine-tuning restricts the potential improvement that the model can achieve. A similar finding can also be observed in the Clinical LLaMA fine-tuning whose overall performance does not differ from the original fine-tuning. This finding is further supported by the improvement in the AUROC scores of LLaMA and PMC-LLaMA with trainable Clinical LLaMA-LoRA, which achieve 75.13% and 72.23% macro-averaged AUROC scores, respectively. These represent substantial improvements from the vanilla fine-tuning performance, 58.61% and 60.51% AUROC scores.
@@
By considering Clinical LLaMA-LoRA as the "delta-updating" outcome of the domain-adaptive pretraining, we can view the downstream fine-tuning process as an additional "delta-updating" step. To investigate the impact of this approach, we conduct experiments by adding a Downstream LLaMA-LoRA to LLaMA and PMC-LLaMA models that were already equipped with Clinical LLaMA-LoRA. From Table 3, we can observe that Downstream LLaMA-LoRA fails to improve the performance of LLaMA and PMC-LLaMA with frozen Clinical LLaMA-LoRA. On the other hand, improvement can be observed when adding Downstream LLaMA-LoRA to LLaMA with trainable Clinical LLaMA-LoRA. This combination of LLaMA with trainable Clinical LLaMA-LoRA and Downstream LLaMA-LoRA achieves the highest macro-averaged AUROC score of 76.07%. The macro-averaged AUROC score of Clinical LLaMA-LoRA was almost similar to that of PMC-LLaMA with LoRA, suggesting similar efficacy between Clinical LLaMA-LoRA and the full fine-tuning process that PMC-LLaMA has undergone. Moreover, Clinical LLaMA-LoRA offers the advantage of reduced computational resources and training time, which is aligned with the requirements of practical implementation in clinical settings.
Overall, our proposed method manages to achieve better performance in comparison to clinically trained models. We also provide a comparison with the state-of-the-art method of PMV, mortality, and length of stay predictions, called BEEP Naik et al. (2022), which leverages retrieval augmentation method to provide more contextual information to the model during inference. The comparison is only partial as BEEP models were not evaluated on the diagnosis and procedure prediction tasks. As shown in Appendix C, our best-performing model achieves a 70.03% averaged AUROC score, which is slightly worse compared to the best-performing BEEP model with 72.26% averaged AUROC score. However, it is worth noting that our proposed method and the state-of-the-art method are complementary to each other. Hence, future work may explore the possibility of combining the two approaches.